OpenStudy (osanseviero):

I have the function sqrt(1+x) How can I find the taylor polynomial in the point zero?

OpenStudy (anonymous):

you can use that binomial stuff, or just grind it til you find it

OpenStudy (anonymous):

\[f(0)=1\\ f'(x)=\frac{1}{2}(1+x)^{-\frac{1}{2}},f'(0)=\frac{1}{2}\] is a start

OpenStudy (osanseviero):

ohh I see so i just create a poynomial that follows the derivatives

OpenStudy (anonymous):

\[c_n=\frac{f^{n}(0)}{n!}\] is the formula to use

OpenStudy (osanseviero):

So that will be like 1 + x/2 -x^2/8...

OpenStudy (anonymous):

it is easiest to find a formula if you do not compute any numbers for example \[f''(x)=-\frac{1}{2}\times \frac{1}{2}(1+x)^{-\frac{3}{2}}\] \[f''(0)=-\frac{1}{2^2}\] \[c_2=\frac{1}{2^2\times 2!}\]

OpenStudy (anonymous):

no i screwed up somewhere, that is wrong

OpenStudy (anonymous):

\[f^{(3)}(x)=\frac{3}{2}\times \frac{1}{2}\times \frac{1}{2}(1+x)^{-\frac{5}{2}}\] that is better i flipped by mistake

OpenStudy (anonymous):

then \[c_3=\frac{3}{2^3\times 3!}=\frac{1}{2^4}\]

OpenStudy (osanseviero):

isnt any of them supposed to be negative?

OpenStudy (osanseviero):

Like the second derivative

OpenStudy (anonymous):

yes it alternates for sure

OpenStudy (anonymous):

if you know how to do that generalized binomial thing (probably in your text book) you can write it as \[\sqrt{x+1}=(1+x)^{\frac{1}{2}}=\sum _{n=0}^{\infty } \binom{\frac{1}{2}}{n} x^n\]

OpenStudy (osanseviero):

ok :) thanks

OpenStudy (anonymous):

yw

OpenStudy (osanseviero):

is the fourth derivative -15/2^4 ?